# Holes in Latent Space: Topological Signatures Under Adversarial Influence

### v1

@@ Abstract
Understanding how adversarial conditions affect language models requires techniques that capture both global structure and local detail within high-dimensional activation spaces. We propose persistent homology (PH), a tool from topological data analysis, to systematically characterize multiscale latent space dynamics in LLMs under two distinct attack modes—backdoor fine-tuning and indirect prompt injection. By analyzing six state-of-the-art LLMs, we show that adversarial conditions consistently compress latent topologies, reducing structural diversity at smaller scales while amplifying dominant features at coarser ones. These topological signatures are statistically robust across layers, architectures, model sizes, and align with the emergence of adversarial effects deeper in the network. To capture finer-grained mechanisms underlying these shifts, we introduce a neuron-level PH framework that quantifies how information flows and transforms within and across layers. Together, our findings demonstrate that PH offers a principled and unifying approach to interpreting representational dynamics in LLMs, particularly under distributional shift.
@@ 1 Introduction
A comprehensive understanding of the latent space of Large Language Models (LLMs) requires a multiscale approach. At the local scale, single neurons and sparse factors encode word pieces, part-of-speech tags, and punctuation cues (tenney2019bertrediscoversclassicalnlp, ; hewitt-manning-2019-structural, ); at an intermediate scale, activation trajectories within middle layers preserve factual associations and maintain discourse coherence (meng2023locatingeditingfactualassociations, ); at the global scale, broader activation regions separate diverse knowledge domains (burns2024discoveringlatentknowledgelanguage, ). However, most existing empirical works assume and focus on linear structure, thus, they overlook potential nonlinear transformations on high-dimensional activations (bruelgabrielsson2020topologylayermachinelearning, ; engels2025languagemodelfeaturesonedimensionally, ). Standard single-scale tools such as linear probes, mechanistic interpretability, and representation engineering offer an incomplete picture.
In this paper, we address this gap by studying LLM hidden states using persistent homology (PH)—a technique from topological data analysis (TDA) that captures the shape and size of data. This information is quantified and encoded in a barcode—a summary statistic that represents the span of multi-scale topological features in data. Barcodes show a clear distinction between normal and adversarial activations; see Figure 1. We conduct an extensive study of the topology of LLM representation spaces and identify topological features captured by PH that are robust across architectures, layers, and training nuances, as well as sufficiently sensitive to identify meaningful alterations due to adversarial intervention. Our key results can be summarized as follows.
@@
  • We model the global layer-wise PH features across six state-of-the-art LLM architectures, and achieve a near-perfect separation of the topology of normal vs. adversarial activations, demonstrating that PH alone can robustly capture the structural deformations induced by adversarial influence.
  • We interpret the PH-derived features to characterize the effect of adversarial attacks on the representation space. Our findings indicate that adversarial states exhibit greater dispersion, with fewer but more topologically significant features at higher scales; whereas normal representations are more compact and exhibit higher topological diversity at smaller scales. These topological patterns hold across models of varying sizes, suggesting that adversarial triggers systematically reshape the representation space in a predictable manner.
  • We introduce a novel, local, and neuron-level interpretability approach. By mapping neuron activations from pairs of layers into 2D coordinates and applying PH, we capture fine-grained structural changes and information flow within the network. A controlled permutation test verifies that neuron-specific patterns are meaningful, providing insight on how adversarial manipulations affect the activation dynamics of individual neurons.
1.1 Related Work
TDA methodology coupled with machine learning has been an active approach to analyzing text data in recent years; see uchendu2024unveilingtopologicalstructurestext for a comprehensive survey of TDA applications in NLP.
TDA for LLM Representations. PH has been used to track layer-wise topology in BERT and RoBERTa (chauhan2022bertopsstudyingbertrepresentations, ; garciacastellanos2024relativerepresentationstopologicalgeometric, ) and to compare models via a persistence-similarity metric (gardinazzi2024persistenttopologicalfeatureslarge, ). It has also been used to detect adversarial image examples provided to a single CNN (gebhart2017adversarydetectionneuralnetworks, ), to detect out-of-distribution text provided to variants of BERT (pollano2024detectingoutofdistributiontextusing, ), and as a proxy for robustness across NLP tasks (chauhan2022bertopsstudyingbertrepresentations, ). Our work is the first to model six state-of-the-art LLMs under two orthogonal attack modes, showing that PH yields architecture-agnostic adversarial signals.
Linear Geometry vs. Higher-Order Structure. Recent work finds that LLMs often store knowledge along simple linear directions that encode semantic facts, world models, and task vectors (vaswani2023attentionneed, ; nanda2023emergentlinearrepresentationsworld, ; marks2024geometrytruthemergentlinear, ; gurnee2024languagemodelsrepresentspace, ). Sparse autoencoder probes provide overlapping features, yet their single linear encoder confines these factors to an affine subspace (cunningham2023sparseautoencodershighlyinterpretable, ; li2024geometryconceptssparseautoencoder, ). Topology layers reveal curved, multi-scale manifolds in vision and graphs (bruelgabrielsson2020topologylayermachinelearning, ), and recent work shows that jailbreaks exploit such nonlinear features (kirch2024featurespromptsjailbreakllms, ). Nonetheless, most defenses still rely on linear probes or coarse clustering (chao2023jailbreaking, ; chwang2024androidsknowtheyredreaming, ; abdelnabi2024trackcatchingllmtask, ).
@@ 2 Background
2.1 Persistent Homology
PH serves as a powerful “topological lens”, providing insights into the multi-scale shape and structural complexity of data beyond conventional linear approaches. In our analysis, for a layer $\ell\in\{1,\ldots,L\}$ in the model, the hidden representations (or activations) of $N$ input examples form a point cloud in $\mathbb{R}^{D}$, where $D$ is the dimension of the representation vectors.
Instead of just treating these LLM activations as distinct points in a high dimensional space, we take a multi-scale topological view by incrementally “thickening” them (see Figure 2). We allow a radius parameter to grow, akin to adjusting thresholds in hierarchical clustering or DBSCAN, to reveal how the activations connect to form structures across different scales. At small radii, the points remain disconnected. As the radius increases and the balls around nearby points intersect, we connect these points to form clusters, but also edges, triangles, and higher-dimensional structures, which represent more complex, global interactions extending beyond standard clustering. This construction is called a Vietoris–Rips filtration.
PH tracks how topological features emerge (are born) and eventually disappear (die) as the connectivity within the data increases. Its main output is a persistence barcode—a visual summary showing their lifetime, i.e., how long each feature exists as the radius increases. In this work, we focus on two types of features: connected components, represented by 0-bars, and loops or cycles, represented by 1-bars. The interpretation of these features as bars is given in Figure 2, where the barcode shows two 1-bars corresponding to the two visible circular structures within the data. Each 0-bar corresponds to a point in the point cloud, and its length reflects how long that point remains disconnected from a cluster, as the radius increases, until it eventually joins one. The very long 0-bar represents the cluster of all points that are obtained at the end of the process and never disappears.
@@ 2.1 Persistent Homology
Figure 2: Left: Vietoris–Rips filtration constructed from a sample of 50 points over 2 circles with noise, at four values of the distance threshold $r\in[0,\infty)$. Right: corresponding persistence barcode for the 0- and 1-bars, with vertical lines corresponding to the thresholds displayed on the left.
[…]
@@ 2.2 Adversarial Influence on LLMs
LLM representation space can be characterized by barcode summaries that are distinct when operating under normal conditions versus when under adversarial influence. We demonstrate the distinguishing power of topological characterizations of representation spaces by studying two systematically different attack modes. Indirect Prompt Injection (XPIA) is a technique where attackers add hidden instructions to the input context provided to the LLM, often via retrieval-augmented content like web pages and documents, to override the user’s original prompt with consequences like data exfiltration greshake2023youvesignedforcompromising ; Rehberger2024 . On the other hand, sandbagging via backdoor fine-tuning is a technique that deliberately suppresses a model’s observable capabilities until a hidden trigger phase is provided which unlocks its full performance greenblatt2024stresstestingcapabilityelicitationpasswordlocked . We study the developer-driven variant of sandbagging, where the model is fine-tuned to underperform vanderweij2024aisandbagginglanguagemodels . In summary, XPIA targets the fundamental inability of current LLMs to distinguish between data and instructions (zverev2025llmsseparateinstructionsdata, ), making it challenging to prevent, while sandbagging targets the model fine-tuning process.
3 Method and Experimental Setup
In this section, we give details on the data we analyze and outline the computational pipelines we implement for the global and local analyses. More details on the latter can be found in Section B.1 and Section C.1.
@@ 3.1 Data and Representations
We study the topological characteristics of two adversarial scenarios—(i) XPIA and (ii) sandbagging—by computing the PH of $\mathbb{R}^{4096}$-dimensional activation vectors from the last token of an input $x$. These vectors encode the model’s aggregated interpretation of the input context zou2023representationengineeringtopdownapproach .
TaskTracker (Clean vs. Poisoned). The TaskTracker abdelnabi2024trackcatchingllmtask test set comprises $>62\text{k}$ examples, where each example represents the activation values for an input $x$ that is either clean or poisoned. Each input combines a user prompt with a retrieved data block. In poisoned cases, the data block contains an injected secondary task that deviates from the original user instruction. These injections span code generation codealpaca , jailbreak prompts, and other prompt injection attacks. We analyze representation data from six instruction tuned LLMs; Phi3-mini-4k (3.8B parameters) and Phi3-medium-128k (14B parameters) (abdin2024phi3technicalreporthighly, ), Mistral 7B (jiang2023mistral7b, ), LLaMA3 8B and 70B (grattafiori2024llama3herdmodels, ), and Mixtral-8x7B (jiang2024mixtralexperts, ).
Sandbagged Models (Locked vs. Elicited). We fine-tuned Mistral 7B and LLaMA3 8B on a corpus of 3,347 synthetic questions inspired by the WMDP benchmark li2024wmdpbenchmarkmeasuringreducing using LoRA hu2021loralowrankadaptationlarge . Following the methodology of vanderweij2024aisandbagginglanguagemodels , we created models with two operational modes: locked (default state) and elicited (activated by a specific password in the prompt). We extracted final-token representations from each layer during inference with evaluation prompts in both states. Further implementation details are provided in Appendix C.3.
@@
3.2 Global Layer-Wise Analysis: Descriptive and Inferential Methods on Barcode Summaries
This analysis aims to highlight the intrinsic differences in the topology of normal versus adversarial activations, facilitating interpretations for the underlying causes of these distinctions. To do this, we implemented the pipeline in Figure 4. We used Ripser++ (bauer_ripser_2021, ; zhang2020gpu, ) to compute barcodes based on Vietoris–Rips filtrations, leveraging subsampling techniques (e.g., chazal2015subsampling ), both to reduce the computational cost of PH and to enable statistically robust inference. For each layer, we drew $K=64$ subsamples of $k=4096$ normal representations; and $K=64$ subsamples of $k=4096$ adversarial representations. Implementation details are given in C.4. Following dashti_2023 , we vectorized these barcodes into 41-dimensional barcode summaries (cf. Section 2.1). To reduce redundancy and prevent overfitting, we removed highly correlated variables, ensuring an efficient and informative representation for a more parsimonious model. We performed PCA and computed canonical correlation analysis (CCA) loadings to investigate feature importance. We also trained a logistic regression and computed Shapley values (lipovetski_2001, ) to evaluate the predictive power of features.
3.3 Local Information Flow Analysis: Study of Consecutive and Non-Consecutive Layers
This analysis aims to characterize how information propagates across network layers by examining changes in neuron activations. Specifically, we projected pairs of $D$-dimensional activation vectors from two distinct layers into a 2-dimensional embedding and computed PH. Each of the $D$ points in the resulting $\mathbb{R}^{2}$ point cloud represents a neuron, where the first coordinate corresponds to its activation in one layer and the second to its activation in the other. We performed this analysis on $n=1000$ such activation pairs from normal and adversarial cases, producing corresponding point clouds. In standard architectures, empirical observations reveal strong correlation patterns within these embeddings (Figure 5(a)). The distribution of activation differences for neurons with the same index across consecutive layers is typically centered around zero (Figure 5(b)), indicating that many activations are preserved with minimal change from layer to layer, consistent with high colinear behavior commonly observed between layers. Conversely, neurons undergoing significant transformations correspond to points that deviate substantially from the identity line $y=x$ in the joint space. We uniquely focus on the structures that emerge from these activation differences between layers, and show these carry meaningful representational information. PH provides a principled and sensitive tool for identifying and quantifying these patterns; see Figure 5(c), illustrating how all points, those on the diagonal as well as those deviating away from it, form cycles leading to persistent 1-bars. We validated the statistical significance of the topological signals captured by our method by constructing a control condition that preserves the distribution of activation values while disrupting neuron-wise correspondence. This was done by randomly permuting neuron indices within normalized activation vectors. We further extended this analysis to non-consecutive layers to investigate whether these topological patterns persist across larger network intervals. See Section C.1 for further methodological details.
@@ 4 Results
4.1 Global Results
We present the results of the analysis in Figure 4 for the instruction-tuned Mistral 7B model for the clean vs. poisoned datasets. Similar results for the instruction-tuned LLaMA3 8B and 70B, Phi3-mini-4k, Phi3-medium-128k, and Mixtral-8x7B can be found in Section B.2. We present analogous results for locked vs. elicited in Section B.3, highlighting differences and similarities among results.
Cross-Correlation Analysis of Barcode Summaries. We studied the cross-correlation matrix of the 41-dimensional barcode summaries obtained from the subsamples. The results in Figure 6 show that a growing block of highly correlated features appears in the layers of the model. The mean death of 0-bars emerges as the first prominent feature, so we retained it as the representative of the topological features in this block in our analyses when pruning highly correlated features. However, we remark that the prominence of this statistic in the results of our analysis does not imply a lack of significance for higher-order topological features (specifically, 1-bars). We observe a strong correlation between statistics of the 0- and 1-bars; mathematically, it is known that the deaths of 0-bars are closely linked to the births of 1-bars (this relationship has been explored in the context of Morse theory; see adler for further discussion). Thus, the mean death of 0-bars inherently captures information on 1-bars as well. In light of the cross-correlation results, we discarded all features that have a correlation higher than a threshold of 0.5 with at least one feature present in the analysis, resulting in the features in Table 4. We refer to this data set as the pruned barcode summaries.

### v2

# The Shape of Adversarial Influence: Characterizing LLM Latent Spaces with Persistent Homology

@@ Abstract
Existing interpretability methods for Large Language Models (LLMs) often fall short by focusing on linear directions or isolated features, overlooking the high-dimensional, nonlinear, and relational geometry within model representations. This study focuses on how adversarial inputs systematically affect the internal representation spaces of LLMs, a topic which remains poorly understood. We propose persistent homology (PH), a tool from topological data analysis, as a principled framework to characterize the multi-scale dynamics within LLM activations. Using PH, we systematically analyze six state-of-the-art models under two distinct adversarial conditions, indirect prompt injection and backdoor fine-tuning, and identify a consistent topological signature of adversarial influence. Across architectures and model sizes, adversarial inputs induce “topological compression”, where the latent space becomes structurally simpler, collapsing from varied, compact, small-scale features into fewer, dominant, and more dispersed large-scale ones. This topological signature is statistically robust across layers, highly discriminative, and provides interpretable insights into how adversarial effects emerge and propagate. By quantifying the shape of activations and neuronal information flow, our architecture-agnostic framework reveals fundamental invariants of representational change, offering a complementary perspective to existing interpretability methods.
@@ 1 Introduction
A comprehensive understanding of the latent space of Large Language Models (LLMs) requires a multiscale approach. LLM representations form a conceptual hierarchy, with local-scale individual neurons encoding simple features such as punctuation (Tenney et al., 2019; Hewitt and Manning, 2019), intermediate-scale circuits forming contextual associations (Meng et al., 2023), and global-scale activation patterns representing more abstract concepts (Burns et al., 2024). However, most empirical work assumes a linear structure, neglecting the complex geometry of these high-dimensional activation spaces (Brüel-Gabrielsson et al., 2020; Engels et al., 2025a). This oversight creates a practical security gap in real-world models, allowing diverse attacks to exploit nonlinear features and bypass the prevalent defenses that rely on linear classifiers (Kirch et al., 2024).
In this paper, we address this gap by studying LLM hidden states using persistent homology (PH)—a technique from topological data analysis (TDA) that captures the multi-scale shape of data (Chazal and Michel, 2021). PH is uniquely suited for this task because it provides a coordinate-free summary of relational geometry that is known to be robust to noise (Cohen-Steiner et al., 2007). Unlike methods that project high-dimensional representations onto lower-dimensional subspaces, PH preserves multi-scale structural information through a filtration, capturing both local clustering patterns and global topological features simultaneously. These properties enable direct and meaningful comparisons of latent space structure across different models, input distributions, and fine-tuning stages. This information is quantified and encoded in a barcode—a summary statistic of the evolution of topological features. As shown in Figure 1, these barcodes elucidate a clear distinction between normal and adversarial activations, motivating our deeper investigation.
@@
Our contributions can be summarized as follows.
  • We present a comprehensive study of six state-of-the-art models under two fundamentally different attack modes revealing that adversarial inputs induce consistent topological behavior within the LLM latent space. Specifically, adversarial inputs cause latent representations to become more dispersed, characterized by fewer but more topologically significant large-scale features. In contrast, normal inputs produce a greater diversity of compact, small-scale structures.
  • We show that this phenomenon holds across models ranging from 7B to 70B parameters, suggesting that adversarial triggers systematically reshape the representation space in a consistent and predictable manner that is independent of specific architectures or training procedures.
  • We introduce a novel, neuron-level PH analysis confirms these geometric shifts at a finer scale, revealing a phase transition in the topological complexity of the information flow.
While standard linear classifiers can also separate normal and adversarial states with high accuracy, our topological framework provides an interpretable, geometric explanation for why this separability exists. These findings establish PH as a powerful complementary tool for interpretability and support the view that the success of linear probes may stem from their approximation of more complex, underlying topological structures (Engels et al., 2025b; Park et al., 2024; Yang et al., 2024).
@@ 2 Background
In this section, we outline persistent homology and the barcode summaries we study; we also provide details on the specific types of adversarial influence we investigate.
2.1 Persistent Homology and Persistence Barcodes
Persistent homology is a powerful methodology to quantify the “shape” and “size” of data, which can be applied to diverse input data types, is robust to noise perturbations, captures higher-order relational information and has an inherently interpretable nature. More precisely, PH captures topological features, e.g., connected components, tunnels and loops, or cavities and bubbles, present at different scales in our data.
For our activation data, i.e., point clouds $X\subset\mathbb{R}^{D}$, with $D$ the hidden dimension of the model (typically, $D=4096$) and where each point is the latent representation of the last token in a prompt; the PH pipeline proceeds as follows. The first step is to construct a dynamic, geometric representation of our point cloud. A classical construction involves the Vietoris–Rips complex, which for a scale parameter $\epsilon>0$ is obtained from the $\epsilon$-neighborhood graph, that is, the graph where we connect any two points at distance less than $\epsilon$. The Vietoris–Rips complex goes beyond the pairwise interactions in the $\epsilon$-neighborhood graph including higher-order relational information, namely, interactions between more than two points at the same time, known as simplices: 0-simplices correspond to points, 1-simplices to edges, 2-simplices to triangles, 3-simplices to tetrahedra, and so on. We add a simplex between a subset of 3 or more points to the Vietoris–Rips complex whenever they are all pairwise connected, for instance, we add a triangle if three points in the point cloud are connected in the $\epsilon$-neighborhood graph. This completes the Vietoris–Rips complex construction. Considering all scale parameters $\epsilon$ at the same time, we obtain the Vietoris–Rips filtration: a growing family of geometric spaces where we connect points and add simplices as the parameter $\epsilon$ grows.
PH then leverages algebraic topology to produce the persistence barcode, a collection of bars capturing how the topological features are formed and disappear in the filtration as the scale parameter $\epsilon$ increases. The barcode is stratified in different dimensions, here we focus in dimensions 0 and 1. Bars in the 0-dimensional barcode (or 0-bars) correspond to connected components: at $\epsilon=0$ there are a many bars in the barcode as points in the data, with bars terminating as point get connected in the $\epsilon$-graph. 1-bars represent loops or cycles in the corresponding Vietoris–Rips complex: a bar starts whenever we have added enough edges to enclose a non-trivial hole, and ends when the addition of triangles covers said hole. Usually, the starting point is called the birth and the ending point the death of the bar. An illustrative example of the PH pipeline in a simple point cloud and the corresponding barcode can be found in Figure 2.
@@ 2.1 Persistent Homology and Persistence Barcodes
Figure 2: Left: Vietoris–Rips filtration constructed from a sample of 50 points over 2 circles with noise, at four values of the distance threshold $\epsilon\in[0,\infty)$. Right: corresponding persistence barcode for the 0- and 1-bars, with vertical lines corresponding to the thresholds displayed on the left.
2.2 Persistent Homology in Machine Learning: Barcode Summaries
Persistence barcodes cannot be directly used as input features in a ML model since they do not reside in a Euclidean space (Turner et al., 2014). We circumvent this issue by studying statistics of vectorized barcodes (Ali et al., 2023)—such as the mean, standard deviation, median, or quartiles—of the empirical distributions of the births, deaths, and persistences (lengths) of the bars in a given barcode. We can also study the empirical distribution of the ratios between births and deaths, which have the advantage of being scale invariant; the number of bars, providing a notion of topological diversity; the total persistence, which is given by the sum of the lengths of all bars in the barcode and captures both the number of topological features and their size; and the persistent entropy (Chintakunta et al., 2015; Rucco et al., 2016) of each barcode, which intuitively measures the heterogeneity within the lengths of the bars in the barcode. In all, for each barcode, we compute a 41-dimensional descriptive feature vector that can be used in machine learning tasks, which we call the barcode summary.
@@ 2.3 Adversarial Influence on LLMs
We study whether a topological analysis of LLM representation spaces can distinguish and provide meaningful insight into normal and adversarial information processing conditions. In order to test the generality of our topological framework, we quantify and interpret the effects of two systematically different attack modes, both of which carry a high security impact in practice.
The two attack modes we study are: Indirect Prompt Injection (XPIA), where attackers embed hidden instructions in retrieved content to override a user’s original prompt (Greshake et al., 2023; Rehberger, 2024); and sandbagging via backdoor fine-tuning, which involves deliberately training a model to suppress its capabilities until a secret trigger is provided (Greenblatt et al., 2024; van der Weij et al., 2024). These techniques target fundamentally distinct vulnerabilities: XPIA exploits the model’s core inability to distinguish data from instructions (Zverev et al., 2025), whereas sandbagging infiltrates the fine-tuning process.
3 Experimental Design
In this section we overview details of the data we study, and outline our global and local studies using persistent homology.
@@ 3.1 Data and Representations
We study two adversarial modes (XPIA and sandbagging) by computing the barcodes of point clouds in $\mathbb{R}^{D}$, where each point corresponds to the latent representation of the last token of a given input, or of a 2D embedding described in Section 3.3. These vectors encode the model’s aggregated interpretation of the input context (Zou et al., 2023).
TaskTracker (Clean vs. Poisoned). The TaskTracker (Abdelnabi et al., 2024) test set comprises $62\text{k}$ examples, where each example contains the activation values for an input that is either clean or poisoned. Each input combines a user prompt with a retrieved data block. In poisoned cases, the data block contains an injected secondary task from several standard safety and security benchmarks, including Beavertrails (Ji et al., 2023), Harmbench (Mazeika et al., 2024), and Jailbreakbench (Chao et al., 2024). These attacks are applied across diverse task prompts, such as coding challenges and question-answering, to create realistic test cases. We analyze representation data from six instruction tuned LLMs; Phi3-mini-4k (3.8B parameters) and Phi3-medium-128k (14B parameters) (Abdin et al., 2024), Mistral 7B (Jiang et al., 2023), LLaMA3 8B and 70B (Grattafiori et al., 2024), and Mixtral-8x7B (Jiang et al., 2024).
Sandbagged Models (Locked vs. Elicited). We fine-tuned Mistral 7B and LLaMA3 8B on a corpus of 3,347 synthetic questions inspired by the WMDP benchmark (Li et al., 2024) using LoRA (Hu et al., 2021). Following the methodology of van der Weij et al. (2024), we created models with two operational modes: locked (default state) and elicited (activated by a specific password in the prompt). We extracted final-token representations from each layer during inference with evaluation prompts in both states. Further implementation details are provided in Appendix C.3.
@@
3.2 Global Layer-Wise Analysis
This analysis establishes and explains a consistent topological distinction between normal and adversarial representations, following the pipeline in Figure 4. We used Ripser++ (Bauer, 2021; Zhang et al., 2020) to compute barcodes, leveraging subsampling techniques, both to reduce the computational cost of PH and to enable statistically robust inference. Subsampling approaches in PH are theoretically grounded, as under mild sampling models, persistence diagrams estimated from point clouds converge to the population diagrams with guaranteed rates (Chazal et al., 2015, 2014). For each model layer, we drew $K=64$ subsamples of $k=4096$ normal representations; and $K=64$ subsamples of $k=4096$ adversarial representations. We vectorized the corresponding barcodes into 41-dimensional barcode summaries (cf. Section 2.2), and performed the analysis in Figure 4, see results and further details in Section 4.1.
3.3 Local Information Flow Analysis
This analysis quantifies neuron-level information flow by tracking topological changes in activation patterns between layers. For each pair of layers $\ell$ and $\ell^{\prime}$, we construct a 2D point cloud from their corresponding $D$-dimensional activation vectors. Each of the $D$ points in this embedding has coordinates $(v_{i}^{\ell},v_{i}^{\ell^{\prime}})$, representing the activation of the $i$th neuron in layer $\ell$ and layer $\ell^{\prime}$, respectively.
The rationale for this embedding is that activations between consecutive layers are empirically highly correlated, causing points to cluster near the identity line $y=x$, as shown in Figure 5(a). Significant transformations in network processing are reflected in neurons whose activations deviate from this line, producing topological structures (e.g., loops) that PH captures and quantifies. We apply this analysis to 1000 clean and 1000 adversarial activation samples to compare the resulting topological signatures, which are presented in Section 4.2.
@@ 4 Results
We now present the implementation results of our proposed analyses to the data described above.
4.1 Global Analysis: The Shape of Adversarial Influence
Our global analysis, as outlined in Figure 4, reveals a consistent and highly discriminative topological signature of adversarial influence across all six LLMs. Specifically, we show that adversarial inputs induce a “topological compression” of the latent space. Here, we present the results of quantifying and interpreting the effect of XPIA on Mistral 7B’s latent space. Results for the other five models are relegated to Appendix B.2. Results for the same models but subjected to the backdoor finetuning attack for sandbagging are given in Appendix B.3.
Cross-Correlation Analysis of Barcode Summaries. In Figure 6, a growing block of highly correlated features appears in the cross-correlation matrix of the 41 features of the barcode summaries. To reduce redundancy and prevent overfitting, we removed highly correlated variables, ensuring an efficient and informative representation for more parsimonious models in subsequent analyses. We discarded all features that have a correlation higher than a threshold of 0.5 with at least one feature present in the analysis, resulting in the features in Table 5. We refer to this data set as the pruned barcode summaries. The first feature appearing in this block of highly correlated features is the mean deaths of the 0-bars (the average of their ending points), which is retained in the pruned barcode summaries as representative of the block. However, we remark that the prominence of this statistic in the results of our analysis does not imply a lack of significance for higher-order topological features (specifically, 1-bars). Empirically, there is a strong correlation between statistics of the 0- and 1-bars in our results; theoretically, it is known that the deaths of 0-bars are closely linked to the births of 1-bars (which has been explored using Morse theory; see Adler and Taylor (2011)).